OpenStudy (anonymous):

In a restaurant, Chad tracked the number of children and adults that came in over an hour. During that time, 16 children and 40 adults came in the restaurant. What is the ratio of adults to children that came into the restaurant? A. 2:7 B. 5:3 C. 5:7 D. 5:2

OpenStudy (welshfella):

40 : 16

OpenStudy (welshfella):

divide 40 and 16 by a common factor

OpenStudy (anonymous):

UMM i suck at math i literally am stupid lol i have no idea how to do that

OpenStudy (anonymous):

and theres like 4 other questions lol

OpenStudy (welshfella):

you need a number which will divide exactly into 40 and 16 to get a smaller ratio. for example we could divide each by 4 this will give 40/4 and 16 / 4 = 10:4 But there is another higher number which will give one of the above options what do you think it is ?

OpenStudy (welshfella):

what number greater than 4 will divide into 40 and 16?

OpenStudy (anonymous):

theres alot of numbers greater than 4 but 6 maybe i dk im just guessing

OpenStudy (welshfella):

8 is the one! 40 / 8 = ? 16 / 8 = ?

OpenStudy (welshfella):

LOl - i misread your post I've told you the answer now.

OpenStudy (anonymous):

lol its ok thanks

OpenStudy (welshfella):

6 wont divide into 40 nor 16 but 8 will.

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

so what is the correct ratio?

OpenStudy (welshfella):

do the 2 divisions

OpenStudy (anonymous):

D ?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

40/8 = 5 16/8 = 2 so its 5:2

OpenStudy (welshfella):

40:16 = 5:2

OpenStudy (anonymous):

ok could you help me with the other 4

OpenStudy (welshfella):

I cant sorry gotta go right now

OpenStudy (anonymous):

In a probability experiment, Craig rolled a six-sided die 56 times. The die landed on a number greater than three 35 times. What is the ratio of rolls greater than three to rolls less than or equal to three? A. 3/5 B. 8/5 C. 5/8 D. 5/3